Triangle DEF is similar to triangle PQR. Side DE measures 3, side EF measures 7 and side FD measures 9. Side PQ measures 5x – 25 and side QR measures 6x + 4. Find the value of x.
Answer #1
DE/PQ = 3/(5x -25) = EF/QR = 7/(6x + 4)

18x + 12 = 35x -175

17x = 187
x = 11

DEF has 1/10 the corresponding side lengths of PQR
answered by: icelocsgirl
Answer #2
d
answered by: Abby Heath